300k temperature change in Celsius.
1
Expert's answer
2020-09-21T08:28:17-0400

Let's denote "t" - temperature in Celsius, "T" - tempetarure in Kelvin.

The connection between these units is "T = t + 273.15" and as following "t = T - 273.15"

If "\Delta T = T_2 - T_1 = 300\; K", then

"\Delta t = t_2 - t_1 = T_2 - 273.15 - (T_1 - 273.15) ="

"= T_2 - 273.15 - T_1 +273.15 = T_2-T_1 = \Delta T".

So, "\Delta t = \Delta T = 300^\circ C".

Answer: "\Delta t= 300^\circ C".